OpenStudy (mew55):

r^2=2csc2theta convert the polar equation to rectangular coordinates

OpenStudy (turingtest):

is that \[r^2=2\csc(2\theta)\]or\[r^2=2\csc^2\theta\]?

OpenStudy (mew55):

the first one \[r ^{2}=2\csc2\]\[r ^{2}=2\csc2\]

OpenStudy (mew55):

its suppose to be a theta in the end but idk its not showing

terenzreignz (terenzreignz):

All you need is the fact that \[x = rcos \theta\]\[y = rsin \theta\]

OpenStudy (mew55):

but dont i need to change the csc into 1/sin?

terenzreignz (terenzreignz):

Well, that too. All I wanted was to stress the importance of those equalities somewhat :)

OpenStudy (turingtest):

yes, and that\[\csc(2\theta)=\frac1{\sin(2\theta)}=\frac{\cancel2}{\cancel2\sin\theta\cos\theta}\]

terenzreignz (terenzreignz):

@TuringTest \[*2\csc(2\theta)=\frac{2}{\sin(2\theta)}=\frac{\cancel2}{\cancel2\sin\theta\cos\theta}\] ;) Just sayin'